Biomed shift takeover phone conference

Date: January 6th 2014, 10h40 by mail

Attendees:

  • Franck Michel (CNRS I3S)
  • Jerome Pansanel (CNRS IPHC)
  • Doan Trung Tung (IFI)

Next conference: January 13th 2014, 09h00.

General information

Remind about best practices:

  • Start by following up on open tickets and verifying solved tickets, before submitting new ones.
  • Before submitting a ticket verify that:
    • Another ticket does not already exist for the same problem
    • The resource is not in downtime and is in production status ⇒ you can use the brand new VAPOR portal to do so
    • The alarm can be reproduced manually
  • For CEs: ignore the alarms in the cases described here.

Takeover report

  • No more problem with the VOMS that has a hectic behavior until Dec. 27th. Seems to be fixed now (tickets 99651 and 99680 closed). To be monitored however.
  • I closed a few old tickets that were solved, and I submitted 2 tickets for SEs that were fixed quickly (100009 and 100010).
  • 2 tickets on hold (94746 and 94778), no changes there.
  • I submitted a few new tickets about failing SEs and CEs, + 4 tickets about CEs publishing wrong data (444444 issue).
  • ticket 99878 : an alarm for a CE although it is no longer supposed to support the VO. In fact this is a problem of data publication in the BDII. I copied the LDAP requests in the ticket, please follow up on this too.

SEs

  • SE se01.dur.scotgrid.ac.uk is still in production although it was to be decommissioned by Dec. 31st (97766). To be followed. Would probably be good to add a comment in the ticket to ask the site about the status.
  • ticket 94778 repurposed: now only a problem of data publication but the SE works.

CEs

Nothing to report.

WMS

Nothing to report.